A chemist needs 100 milliliters of a 63% solution but has only 55% and 95% solutions available. Find how many milliliters of each that should be mixed to get the desired solution.
Answer by stanbon(57983)

A chemist needs 100 milliliters of a 63% solution but has only 55% and 95% solutions available. Find how many milliliters of each that should be mixed to get the desired solution.
---------------
Let amount of 55% solution be "x" ml ; Amt of active ingredient = 0.55x ml
---------------
Amount of 95% solution is "100-x" ml ; Amt of active ingre. = 0.95(100-x)
= 95-0.95x ml
---------------------
Amount of 63% solution is 100 ml ; Amt of active ingre. 0.63(100) = 63 ml
---------------
EQUATION:
active + active = active
0.55x + 95-0.95x = 63
-0.40x = -32
x= 80 ml (amount of 55% solution needed)
100-80 = 20 ml (amount of 95% solution needed)
===========
Cheers,
Stan H.
San